Solve each equation
3/x-2 + 1/x+2 =12/x2 - 4
I know that I need to find a common denominator and to reduce the x squared would be:
3/X-2 + 1/X+2 = 12/(x+2)(x-2)

This is where I get confused.
The common denominator is x + 2
3 (x+2) + 1 (x-2) = 12
3x + 6 + 1x - 2 = 12
4x - 4 - 12 = 0
(2x - 3) (2x +4) = 0

Answer by stanbon(75887)

Solve each equation
3/(x-2) + 1/(x+2) =12/(x^2 - 4)
----------------------------------------
Multiply thru by (x^2-4) which is (x-2)(x+2)
----------------------------------------
3(x+2) + (x-2) = 12
3x+6 + x-2 = 12
4x+4 = 12
x+1 = 3
x = 2
----
But x cannot be 2 because of the denominator x-2
=====================
Answer: No solution
================================
Cheers,
Stan H.